OpenStudy (anonymous):

Evaluate. Please show all of your work. -y – (-2y – 4), for y = -3

OpenStudy (anonymous):

plug y=-3 IN PARENTHESIS in

OpenStudy (anonymous):

make sure its inparenthesis

OpenStudy (anonymous):

(-3)-(-2y - 4)

OpenStudy (anonymous):

See thats why in parenthesis the first -3 has another negative outside of it. And the second y should be replaced too

OpenStudy (anonymous):

to separate the subtraction and the negative

OpenStudy (anonymous):

yes so plug it in and show me what you got

OpenStudy (anonymous):

(-3)-(-2(-3)-4)

OpenStudy (anonymous):

There is still the extra negative missing outside the first -3

OpenStudy (anonymous):

do you see why?

razor99 (razor99):

hey i have answered your question

razor99 (razor99):

@Hollywoodz

razor99 (razor99):

hello u there

OpenStudy (anonymous):

he is viewing something else

OpenStudy (anonymous):

ok i see it

razor99 (razor99):

hope it helped :p

OpenStudy (anonymous):

ok so you got it?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

well from where you were with me you just simply everything

OpenStudy (anonymous):

which would end up as 3+24

OpenStudy (anonymous):

I think... maybe this is how it works: -y - (-2y-4) distribute first the - sign, -y +2y +4 then, when y=-3 -(-3) +2(-3) +4 =3 - 6 +4 which is = to 1?